# SN 2021foa: Deriving a continuity between SN IIn and SN Ibn

### v1

@@ 7.1 Physical scenario and asymmetry:
The H profile of SN 2021foa shows a deficit in the flux in the red side of the wing and a systematic blueshift in the line centers from km s-1 to 377 km s-1. This deficit is visible as an overall blueshift of the line after day 40 and this increases with time.
@@
To investigate the case of SN 2021foa, we see that the blueshift increases with time. In the case of a blueshift caused by the radiative acceleration of the CSM, the blueshift should decrease as the luminosity drops, and we expect no significant wavelength dependence for this case. Additionally, in the case of a blueshift caused by the radiative acceleration scenario, the original narrow line photon source should be blueshifted as well which is not true for SN 2021foa (Dessart et al., 2015). For obscuration by the continuum photosphere, the blueshift should be strongest in early times and decrease later on as continuum optical depth drops. For a lopsided CSM structure as well, blueshift should be present from early times which is not consistent with observations of SN 2021foa. So, we propose that the arising blueshift is most-likely arising due to newly formed dust grains in the ejecta or post-shock CSM. We do see a late-time flattening in the optical light curves of SN 2021foa. However, we should remark that other signatures of dust formation which involves increase in NIR flux is not seen in the case of SN 2021foa due to lack of observations and also because it went behind the sun. Nonetheless, dust formation is very common in SNe IIn/Ibn 2006tf (Smith et al., 2008b), 2006jc (Smith et al., 2008a), 2010jl (Maeda et al., 2013) and should be the most-likely case happening for SN 2021foa as shown in the recent paper by Farias (2024).
@@
An interesting point and open question to the behaviour of the SN is if the intermediate/narrow lines are from the CSM, what is going on behind the SNe IIn to SNe Ibn transition is indeed hard to explain. If this would simply reflect the CSM composition, it requires that the fraction of He in the mass-loss wind/ejecta is decreasing toward the SN, the opposite from the standard picture. At the same time, it might be difficult to explain this transition in terms of ionization, as the SN luminosity decreases. So, even though the claims of a single star exploding while transitioning from a LBV to WR phase is explained in past studies, the He envelope in the CSM could be due to a star in binary composition giving rise to narrow emission lines. A detailed theoretical interpretation is essential to describe the plausible scenario giving rise to these kind of SNe.
8 Summary:
@@
We propose that the shoulder in the lightcurve arises due to the geometry of the CSM and the late-time flattening is most-likely arising due to either formation of new dust or pre-existing dust as seen from the systematic blueshift in the H profile at late phases which is confirmed by the presence of thermal dust by Farias (2024).
@@ Acknowledgements
This work makes used of data from the Las Cumbres Observatory global telescope network. The LCO group is supported by NSF grants AST-1911151 and AST-1911225. C.P. acknowledges support from ADAP program grant No. 80NSSC24K0180 and from NSF grant AST-2206657.

### v3

@@ 7.1 Physical scenario and asymmetry
The H profile of SN 2021foa at late phases shows a deficit in the flux in the red side of the wing and a systematic blueshift in the line centers from km s-1 to 377 km s-1. This deficit is visible as an overall blueshift of the line after day 40 and this increases with time.
@@
To investigate the case of SN 2021foa, we see that the blueshift increases with time. In the case of a blueshift caused by the radiative acceleration of the CSM, the blueshift should decrease as the luminosity drops, and we expect no significant wavelength dependence for this case. Additionally, in the case of a blueshift caused by the radiative acceleration scenario, the original narrow line photon source should be blueshifted as well which is not true for SN 2021foa (Dessart et al., 2015). For obscuration by the continuum photosphere, the blueshift should be strongest in early times and decrease later on as continuum optical depth drops. For a lopsided CSM structure as well, blueshift should be present from early times which is not consistent with observations of SN 2021foa. Farias (2024) in their observations ruled out the scenario for the new dust formation in the CDS because for that case, the blueshift of the emission line profiles and red-blue asymmetry exhibits a measurable wavelength dependence, with bluer emission lines exhibiting bluer redshifts than the redder wavelengths. This is not observed both in our and their spectral evolution. Additionally, ejecta dust formation occurs typically 1 yr post explosion reaching temperatures of 1500-2000 K which is also not seen in our spectral evolution. So, we propose that the arising blueshift is most-likely arising due to dust or due to occultation by the photosphere. This dust is most-likely pre-existing at some distance and not newly formed. We do see a late-time flattening in the optical light curves of SN 2021foa. However, we should remark that other signatures of dust formation which involves increase in NIR flux is not seen in the case of SN 2021foa due to lack of observations and also because it went behind the sun. Nonetheless, dust formation is very common in SNe IIn/Ibn 2006tf (Smith et al., 2008b), 2006jc (Smith et al., 2008a), 2010jl (Maeda et al., 2013) and could be a probable case for SN 2021foa.
@@
An interesting point and open question to the behaviour of the SN is if the intermediate/narrow lines are from the CSM, what is going on behind the SNe IIn to SNe Ibn transition is indeed hard to explain. If this would simply reflect the CSM composition, it requires that the fraction of He in the mass-loss wind/ejecta is decreasing toward the SN, which is in accordance with the standard picture. Basically, it could be just ionisation effect and composition effect. Further modelling is required to say more about this. Even though the claims of a single star exploding while transitioning from a LBV to WR phase is explained in past studies, the He envelope in the CSM could also be due to a star in binary composition giving rise to narrow emission lines. A detailed theoretical interpretation is essential to describe the plausible scenario giving rise to these kind of SNe.
8 Summary
@@
We propose that the shoulder in the lightcurve arises due to the geometry of the CSM and the late-time flattening is most-likely arising from the dust or occulation (Farias, 2024) as seen from the systematic blueshift in the H profile at late phases. The dust also is pre-existing at some distance.
@@ Acknowledgements
We thank the anonymous referee for their valuable comments/suggestions which have significantly helped in improving the manuscript. This work makes used of data from the Las Cumbres Observatory global telescope network. The LCO group is supported by NSF grants AST-1911151 and AST-1911225. C. Pellegrino acknowledges support from ADAP program grant No. 80NSSC24K0180 and from NSF grant AST-2206657. ND and KM acknowledge the support from BRICS grant DST/ICD/BRICS/Call-5/CoNMuTraMO/2023 (G) funded by the Department of Science and Technology (DST), India. MS acknowledges the financial support provided under the National Post Doctoral Fellowship (N-PDF; File Number: PDF/2023/002244) by the Science & Engineering Research Board (SERB), Anusandhan National Research Foundation (ANRF), Government of India. K. Maeda acknowledges the JSPS Kakenhi grant (JP24H01810, 24KK0070). S. Schulze is partially supported by LBNL Subcontract 7707915. We thank the staff of the GMRT that made these observations possible. GMRT is run by the National Centre for Radio Astrophysics of the Tata Institute of Fundamental Research. The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc.
@@
The data presented in this paper will be provided upon request. All the spectra will be made publicly available in WiseRep and Zenodo.